OpenStudy (anonymous):

Find the standard form of the equation of the parabola with a focus at (3, 0) and a directrix at x = -3

OpenStudy (amistre64):

what ahve you been able to process so far?

OpenStudy (amistre64):



OpenStudy (amistre64):

looks to be a x = ay^2 format, and i can never really recall the parabolic equations :) so I just equate: x = ay^2 , with y = 2x, when x=3 3 = a(2*3)^2 a = 3/36